OpenStudy (anonymous):

the value of √-9 . √-16

OpenStudy (anonymous):

what's the problem?

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{-9}\sqrt{-16}\] \[\large \sqrt{-1*9}\sqrt{-1*16}\] \[\large \sqrt{-1}*\sqrt{9}\sqrt{-1}*\sqrt{16}\] \[\large \sqrt{-1}*\sqrt{-1}*\sqrt{9}*\sqrt{16}\] \[\large i*i*\sqrt{9}*\sqrt{16}\] See what to do from here?

OpenStudy (anonymous):

\[\sqrt{-1}=i\]

OpenStudy (anonymous):

it says the answer is -12

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

nvm i get it! ty!

jimthompson5910 (jim_thompson5910):

oh ok, you sure?

OpenStudy (anonymous):

yea thanks! :D

jimthompson5910 (jim_thompson5910):

you're welcome